how do I graph 3y^2-x^2+12y+2x=5
Answer by stanbon(75887)

how do I graph 3y^2-x^2+12y+2x=5
---------------
You should convert the equation to a standard form
by completing the square on the y-terms and on the
x-terms separately:
------------
3(y^2+4y+4)-(x^2-2x+1) = -5+12-1
3(y+2)^2 - (x-1)^2 = 6
3(y+2)^2/6 - (x-1)^2/6 = 1
(y+2)^2/2 - (x-1)^2/6 = 1
-----------------------
The graph is a hyperbola with center at (1,-2)
The hyperbola opens up and down.
The y -intercept can be found by letting x=0 and
finding two values for "y".
================================
Cheers,
Stan H.